OpenStudy (anonymous):

how do i find the minimum acceleration attained on the interval 0<=t<=4 by the particale whos velocity is given by v(t)=t^2-4t^2-3t+2

OpenStudy (bahrom7893):

differentiate the acceleration

OpenStudy (bahrom7893):

i think u meant v(t)=t^3-4t^2-3t+2

OpenStudy (bahrom7893):

a=v'=3t^2-8t-3 a'=6t-8=0 6t=8 t=8/6

OpenStudy (bahrom7893):

test that and the endpoints

OpenStudy (bahrom7893):

Plug in 0, 8/6, and 4 into 3t^2-8t-3 and see where you get the smallest value.

OpenStudy (bahrom7893):

0 should be the min acceleration.. i think

OpenStudy (experimentx):

differentiate it, v' = 3*t^2 - 8*t - 3 find critical points, v' = 0 =? t = 3*t^2 - 8*t - 3 => t =3, -1/3 v '' = 6*t - 8 put the values of t at v'', where you get negative value you get maximun ie at t=1/3

OpenStudy (bahrom7893):

@ExperimentX he's asking for the MINIMUM of ACCELERATION

OpenStudy (bahrom7893):

you gave him the MAXIMUM of VELOCITY

OpenStudy (experimentx):

Oo mistake .. for minimun you get at t = 3, where v'' has negative value

OpenStudy (experimentx):

haha .. .plus it was outside boundary

OpenStudy (bahrom7893):

experiment stop confusing ppl lol

OpenStudy (bahrom7893):

Okay here: You are given velocity, you need to find the min acceleration. Thus you will have to differentiate acceleration and find where its minimum occurs. But to get the formula of acceleration you must differentiate velocity. Got it?

OpenStudy (bahrom7893):

So when u're differentiating acceleration, you're actually taking the 2nd derivative of velocity.

OpenStudy (experimentx):

Oops sorry, i thought i should give the minimum value of velocity. okay that would be minimum value of v' right, that would be 8/6

OpenStudy (bahrom7893):

no, because u have to test endpoints too

OpenStudy (bahrom7893):

plug them all in into acceleration equation.